OpenStudy (anonymous):

15. Find the sum of all multiples of 3 between 100 and 1000000.?

OpenStudy (anonymous):

333300

OpenStudy (anonymous):

(1000000-100)/3

OpenStudy (anonymous):

Σ[1,333300] (3k + 99) = (333300/2) * (102 + 999999) = 166,666,831,650